OpenStudy (anonymous):

Convert the equations you arrived at in question 2 and question 4 into slope-intercept form. 2: y-4=1/2(x-1): (I think it would be y=-1x+8) 4: y-5=1/2(x-3):(Not sure)

OpenStudy (anonymous):

2(y-4)=(x-1) 2y-8=x-1 2y-x=-1+8 2y=x+7

OpenStudy (anonymous):

2(y-5)=(x-3) 2y-10=x-3 2y=x-3+10 2y=x+7

OpenStudy (anonymous):

both comes out to be same..lol

OpenStudy (anonymous):

Oh wow! Thanks for writing it out like that! It helps a lot! ^-^

OpenStudy (anonymous):

welcome!! i hope u undrstood :-)

OpenStudy (anonymous):

I do know! Thanks